Circular No. 8015

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hx IN PGC 23727 J. Graham and W. Li, University of California at Berkeley, report the

LOTOSS (cf. IAUC 7906) discovery of an apparent supernova (mag ∼ 18.3) on unfiltered KAIT images taken on Nov. 11.5 and 12.5 UT. SN 2002hx is located at α = 8^{h}27^{m}39^{s}.43, δ = −14^{o}47′15′′.7 (equinox 2000.0), which is 4′′.2 west and 20′′.3 south of the nucleus of PGC 23727. A KAIT image taken on Oct. 26.5 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2002hw IN UGC 52 M. Hamuy, Carnegie Observatories, writes that a spectrum (range 320–

930 nm) of SN 2002hw (cf. IAUC 8014), obtained on Nov. 11.05 UT with the Las Campanas Baade 6.5-m telescope (+ Boller & Chivens spectrograph), shows it to be a type-Ia supernova near maximum light. Adopting the NED redshift of 5257 km/s, the minimum of the Si II 633.5-nm absorption yields an expansion velocity of 11500 km/s. Na I D absorption at the redshift of the host galaxy is observed with an equivalent width of 0.26 nm, which suggests significant reddening.

A. V. Filippenko and R. Chornock, University of California, Berkeley, communicate: “Inspection of CCD spectra (range 310–1000 nm) obtained on Nov. 11 UT with the Keck I 10-m telescope (+ LRIS) shows that the object is of type Ia, close to maximum brightness. Relative to the Si II 615-nm absorption line, the Si II/Ti II 580-nm absorption is stronger than normal, suggesting that the object might be somewhat subluminous.”

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002hw, obtained by P. Berlind on Nov. 11.19 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a reddened type-Ia supernova before maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 5257 km/s for the host galaxy, the supernova expansion velocity is ∼ 11000 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 4 ± 2 days before maximum light. The spectrum shows a relatively strong Na I D absorption at the velocity of the host galaxy (equivalent width 0.2 nm). The dust maps of Schlegel et al. (1998 Ap.J. 500, 525) indicate a Galactic reddening of E(B − V) = 0.11. In addition, the overall shape of the spectrum is flat toward the blue, implying that the supernova is extinguished by dust.

2002 November 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green